A Client has decided to use a Cost Plus Incentive Fee contract price arrangement to motivate the Contractor to be more efficient and productive in executing the project. The Contractor’s Fee will be calculated based on a Sliding Fee approach by considering a $110,000,000 target price and a 4% Base (R).

Calculate Contractor’s Fee and Total Payment to the Contractor for two following scenarios:

  • Scenario 1) Actual Cost of Project = $102,000,000
  • Scenario 2) Actual Cost of Project = $124,000,000
Answer #1

Given data :-

Target price = $110,000,000

Base (R) = 4%

Scenario 1). :-

Actual cost of project = $102,000,000

Contractor fee :--

Now , formula for calculating sliding fee by sliding approach is

Sliding fee = R (T - A)

Where , R= base Percentage value

T = target price

A= actual cost of project

Sliding fee = 4% ×(110,000,000 - 102,000,000)

Sliding fee = (4/100) ×( 80,000,000)

fee = $ 320,000C

Contractor fee =(actual cost of project)+(sliding fee)

Contractor fee= (102,000,000 + 320,000)

Contractor fee= $ 102,320,000

Total payment of contractor:-T

Total payment of contractor is

= (contractor fee)+(overheads percentage× contractor fee)

Where overhead percentage includes contactor profit for their work which is between 10 to 20 %

Taking overhead percentage= 20%

Now, total payment of contractor is

= (102,320,000) + ( 20% × 102,320,000)

Total payment of contractor = $ 122,784,000

Scenario 2) :--

Actual cost of project= $124,000,000

Contractor fee :-

   Sliding fee = R × (T - A)

Sliding fee = 4% ×( 110,000,000 - 124,000,000)

= (4/100) ×( -14,000,000)

Sliding fee= - $560,000

this means the contractor is penalized by the amount of $ 560,000

Contractor fee= (124,000,000 - 560,000)

Contractor fee= $123,440,000

Total payment of contractor is:-

Total payment of contractor

= Contractor fee + (overhead× contractor fee)

= 123,440,000 + ( 20% × 123,440,000)

Total payment of contractor= $148,128,000